OpenStudy (anonymous):

x^3(sinx+1) Differentiate!

OpenStudy (anonymous):

\[x^3(sinx+1)\]

OpenStudy (campbell_st):

product rule u = x^3 du = 3x^2 v = sin(x) dv = cos(x) dy/dx = x^3cos(x) + 3x^2sin(x) = x^2(xcos(x) + sin(x))

OpenStudy (campbell_st):

oops v = sin(x) + 1 dv = cos(x) dy/dx = x^3cos(x) + 3x^2(sin(x)+1)

OpenStudy (anonymous):

Thanks :)